### v1

@@ 5.2 Quadratic program and linear program
$p_{\bm{x}_{1}}(\bm{\lambda})$ $:=\prod_{j\in[p]}\frac{\lambda_{j}^{x_{1,j}}\exp(-\lambda_{j})}{x_{1,j}!}=\Big{(}\prod\limits_{j\in[p]}x_{1,j}!\Big{)}^{-1}\exp\Big{(}\underbrace{\big{(}\ln\lambda_{1},...,\ln\lambda_{p}\big{)}^{\top}}_{\bm{\eta}(\bm{\lambda})^{\top}}\underbrace{(x_{1,1},...,x_{1,p})}_{\bm{\gamma}(\bm{x}_{1})}-\underbrace{\bm{1}^{\top}\bm{\lambda}}_{=A(\bm{\lambda})}\Big{)}$
@@
$p_{\bm{x}_{2}}(\bm{\theta})$ $:=\frac{1}{(2\pi)^{\frac{p}{2}}\sqrt{|\bm{\Sigma}|}}\exp\Big{(}-\frac{1}{2}(\bm{x}_{2}-\bm{\theta})^{\top}\bm{\Sigma}^{-1}(\bm{x}_{2}-\bm{\theta})\Big{)},$
where $\bm{a}_{j}$ and $\bm{q}_{j}$ are the $j$-th column of $\bm{A}$ and $\bm{Q}$, respectively, for $j\in[p]$, namely, $\bm{A}=(\bm{a}_{1},\ldots,\bm{a}_{p})$ and $\bm{Q}=(\bm{q}_{1},\ldots,\bm{q}_{p})$. It is easy to see that (65) is a member of the exponential family.
@@
$G_{1}(\bm{x}_{1})=1,\quad G_{2}(\bm{x}_{2}):=\sqrt{\frac{|\bm{\Sigma}|}{|\tilde{\bm{Q}}+\bm{\Sigma}^{-1}|}}\exp\Big{(}-\frac{1}{2}\bm{x}_{2}^{\top}\tilde{\bm{Q}}\bm{x}_{2}-\tilde{\bm{b}}^{\top}\bm{x}_{2}\Big{)},$
$\tilde{\bm{Q}}:=\bm{\Sigma}^{-1}(\bm{\Sigma}^{-1}-\hat{\bm{Q}})^{-1}\bm{\Sigma}^{-1}-\bm{\Sigma}^{-1},\quad\mbox{and}\quad\tilde{\bm{b}}:=-\bm{\Sigma}^{-1}(\bm{\Sigma}^{-1}-\hat{\bm{Q}})^{-1}\bm{b},$
the condition (54) for (64) is fulfilled with $p_{\bm{x}}(\bm{\theta})$ defined by (65), $f(\bm{\theta})=\frac{1}{2}\bm{\theta}^{\top}\bm{Q}\bm{\theta}+\bm{b}^{\top}\bm{\theta}$, and $\Theta=\mathbb{R}^{p}_{+}$.
@@
For $(\bm{\theta},\bar{\bm{\theta}})\in\mathrm{int}\Theta\times\mathrm{int}\Theta$, the surrogate $u(\bm{\theta}|\bar{\bm{\theta}})$ is a convex function of the form:
@@
and the EM algorithm for (64) is described in Algorithm 6.
Algorithm 6 EM algorithm for quadratic program (64)
1:Let $\bm{\theta}^{(0)}\in\mathbb{R}^{p}$ be an interior point of the feasible region of (64), and set $t\leftarrow 1$.
@@
$\bm{\theta}^{(t)}\leftarrow\arg\min_{\bm{\theta}\in\mathbb{R}^{p}}\left\{\begin{array}[]{r}-\sum\limits_{j\in[p]}\Big{(}\bm{a}_{j}^{\top}\bm{\theta}+\mu^{(t-1)}_{j}\ln(c_{j}-\bm{a}_{j}^{\top}\bm{\theta})\Big{)}+\frac{1}{2}\bm{\theta}^{\top}\bm{\Sigma}^{-1}\bm{\theta}+(\bm{\nu}^{(t-1)})^{\top}\bm{\theta}\end{array}\right\},$ (67)
@@
Suppose that $\bm{Q}\in\mathbb{S}^{p}_{+}$, $\bm{A}$ is of full rank, and (64) has an interior point and optimum. Then the sequence, $\{\bm{\theta}^{(t)}\}_{t\geq 0}$, generated by Algorithm 6 converges to an optimal solution $\bm{\theta}^{\star}$ of (64). Moreover, if $X^{\star}:=\{\bm{\theta}|f(\bm{\theta})=f(\bm{\theta}^{\star})\}\neq\emptyset$, the sequence $\{\bm{\theta}^{(t)}\}_{t\geq 0}$ converges to an optimal solution of (64).
@@
$\begin{array}[]{|ll}\underset{\bm{\theta}\in\mathbb{R}^{p}}{\mbox{maximize}}&\frac{1}{2}\bm{\theta}^{\top}\bm{Q}\bm{\theta}+\bm{b}^{\top}\bm{\theta}\\ \mbox{subject to}&\bm{l}\leq\bm{\theta}\leq\bm{u},\\ \end{array}$ (70)
which is of the form (64) with $\bm{c}=(\bm{u}^{\top},-\bm{l}^{\top})^{\top}$ and $\bm{A}=(\bm{I},~{}-\bm{I})$. For $\bm{\theta}^{(t-1)}=\bar{\bm{\theta}}$, the $\bm{\theta}^{(t)}$ is obtained as the solution $\bm{\theta}$ of
@@
If $\bm{\Sigma}$ is a diagonal matrix $\mathrm{diag}(\bm{\sigma})$ for $\sigma_{j}>0~{}(j\in[p])$, the $j$-th component of $\bm{\theta}^{(t)}$ is obtained as a solution satisfying $\theta_{j}\in(l_{j},u_{j})$ and
$-\frac{\bar{\theta}_{j}-l_{j}}{\theta_{j}-l_{j}}+\frac{u_{j}-\bar{\theta}_{j}}{u_{j}-\theta_{j}}-\kappa_{j}+\sigma_{j}^{-1}\theta_{j}=0,$ (71)
where $\bar{\theta}_{j}=\theta^{(t-1)}_{j}$ and $\kappa_{j}:=[(\bm{\Sigma}^{-1}-\bm{Q})\bar{\bm{\theta}}-\bm{b}]_{j},~{}j\in[p]$. Note that (71) has a unique solution since the left-hand side increases monotonically from $\infty$- ∞ to $\infty$ on the interval $(l_{j},u_{j})$, and it is easy to see that the solution is obtained via the explicit solution formula for the cubic equation:
@@
which is obtained by multiplying (71) by $-\sigma_{j}(\theta_{j}-l_{j})(u_{j}-\theta_{j})$. Accordingly, the MM algorithm results in a simple component-wise formula like (35) and (43). (70) can also be approached using (35), but it is interesting to note that this one is different from (35).
@@
We can construct another EM/MM algorithm based on the combination of the Poisson and normal distributions for another form of QP, which is explained in Appendix B since the construction is parallel to the one for (64). On the other hand, using this derivation, the EM/MM algorithm for LP can be constructed based only on the Poisson and does not require the normal distribution. See the appendix for the details.
6 Concluding remarks
@@
In particular, while we collectively present several linkages between optimization problems and probability distributions, some problems can be approached via different reformulations and distributions. For example, the quadratic minimization over the rectangle can be solved not only via the approach (ii) in Table 1 but also (iv) and (vi), as demonstrated in the final paragraph in the previous section. Further comparison among different formulations or that among different distributions are also to be investigated.
Relations to information geometry may be an interesting research topic as well, since its relation to interior point algorithms has been occasionally pointed out in the literature [30, 26, 17], but the analysis based on the exponential family distributions is not examined yet except in the context of statisitical EM algorithms (e.g., [2, 7]).
@@
Acknowledgements Research of the second author is supported in part by MEXT Grant-in-Aid 20H00285 and 24K01113.
Appendix A Derivation of formulae and proof of propositions
A.1 Proof of Proposition 3.1.
(a):
The claim is confirmed by simply substituting $\tilde{A}(\bm{\theta}^{(t-1)})$ and $\tilde{\beta}(\bm{x})$ into the formula to be proved.
(b):
By the definition (5) of the surrogate function $u$, we have
$u(\bm{\theta}|\bar{\bm{\theta}})$ $=-\mathbb{E}_{q_{\bm{x}}(\bar{\bm{\theta}})}[\ln G(\bm{X})]-\mathbb{E}_{q_{\bm{x}}(\bar{\bm{\theta}})}[\ln p_{\bm{X}}(\bm{\theta})]+\mathbb{E}_{q_{\bm{x}}(\bar{\bm{\theta}})}[\ln q_{\bm{X}}(\bar{\bm{\theta}})]$
@@ Appendix A Derivation of formulae and proof of propositions
$=-\mathbb{E}_{q_{\bm{x}}(\bar{\bm{\theta}})}[\ln p_{\bm{X}}(\bm{\theta})]+\mathbb{E}_{q_{\bm{x}}(\bar{\bm{\theta}})}[\ln p_{\bm{X}}(\bar{\bm{\theta}})]+f(\bar{\bm{\theta}})$
$=-\mathbb{E}_{q_{\bm{x}}(\bar{\bm{\theta}})}[\bm{\gamma}(\bm{X})]^{\top}\big{(}\bm{\eta}(\bm{\theta})-\bm{\eta}(\bar{\bm{\theta}})\big{)}+A(\bm{\theta})-A(\bar{\bm{\theta}})+f(\bar{\bm{\theta}}),$ (14)
which shows the formula (14).
(c):
Although the statement (c) is known in the context of EM algorithms, we here present its proof to make the article self-contained. Note
$\nabla A(\bm{\theta})$ $=\frac{1}{\exp(A(\bm{\theta}))}\nabla\Big{(}\int_{\bm{x}\in\Xi}\beta(\bm{x})\exp\big{(}\bm{\eta}(\bm{\theta})^{\top}\bm{\gamma}(\bm{x})\big{)}\mathrm{d}\bm{x}\Big{)}$ (($)$ )
@@
$=\nabla\bm{\eta}(\bm{\theta})^{\top}\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}[\bm{\gamma}(\bm{X})],$ (72)
where all the gradient operations $\nabla$ apply with respect to $\bm{\theta}$. Owing to (14) and (72), we have
$\nabla_{\bm{\theta}}u(\bm{\theta}|\bm{\theta}^{(t-1)})$ $=\nabla A(\bm{\theta})-\nabla\bm{\eta}(\bm{\theta})^{\top}\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]$ (($)$ )
$=\nabla\bm{\eta}(\bm{\theta})^{\top}\big{(}\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}[\bm{\gamma}(\bm{X})]-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]\big{)},$ (($)$ )
@@
(d):
Observe first that (72) holds true also for $\tilde{A}(\bm{\theta})$ and $q_{\bm{x}}(\bm{\theta})$ in place of $A(\bm{\theta})$ and $p_{\bm{x}}(\bm{\theta})$, respectively, namely,
$\nabla\tilde{A}(\bm{\theta})=\nabla\bm{\eta}(\bm{\theta})^{\top}\mathbb{E}_{q_{\bm{x}}(\bm{\theta})}[\bm{\gamma}(\bm{X})].$ (73)
@@
which leads to (15) by multiplying the inverse of $\nabla\bm{\eta}(\bm{\theta})^{\top}$.
A.2 Proof of Corollary 3.2
@@
which equals 00 at $\bm{\theta}=\bar{\bm{\theta}}$.
A.3 Proof of Theorem 3.3
@@
$\nabla\ln p_{\bm{x}}(\bm{\theta})$ $=\nabla\left(\ln\beta(\bm{x})-A(\bm{\theta})+\bm{\eta}(\bm{\theta})^{\top}\bm{\gamma}(\bm{x})\right)$ (($)$ )
@@
$=\nabla\bm{\eta}(\bm{\theta})^{\top}\bm{\gamma}(\bm{x})-\nabla\bm{\eta}(\bm{\theta})^{\top}\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}[\bm{\gamma}(\bm{X})]$ (($)$ )
@@
$=\bm{I}$ (75)
since we have $\int_{\bm{x}\in\Xi}\bm{\Omega}\bm{\gamma}(\bm{x})\nabla p_{\bm{x}}(\bm{\theta})^{\top}\,\mathrm{d}\bm{x}=\bm{I}$ and $\int_{\bm{x}\in\Xi}\nabla p_{\bm{x}}(\bm{\theta})\,\mathrm{d}\bm{x}=\bm{0}$ by differentiating (17) and $\int_{\bm{x}\in\Xi}p_{\bm{x}}(\bm{\theta})\,\mathrm{d}\bm{x}=1$ with respect to $\bm{\theta}$, respectively. Therefore, we have
@@
$=\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}[\left(\bm{\Omega}\bm{\gamma}(\bm{X})-\bm{\theta}\right)\left(\bm{\Omega}\bm{\gamma}(\bm{X})-\bm{\theta}\right)^{\top}]\big{(}\nabla\bm{\eta}(\bm{\theta})^{\top}\bm{\Omega}^{-1}\big{)}^{\top}$ (($)$ )
@@
so we have $\mathbb{V}_{p_{\bm{x}}(\bm{\theta})}[\bm{\Omega}\bm{\gamma}(\bm{X})]=\big{(}\nabla\bm{\eta}(\bm{\theta})^{\top}\bm{\Omega}^{-1}\big{)}^{-1}$, establishing the second equality of (18). With this, the first equality of (18) can also be confirmed as
$\bm{I}_{p_{\bm{x}}}(\bm{\theta})$ $=\nabla\bm{\eta}(\bm{\theta})^{\top}\bm{\Omega}^{-1}\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}\big{[}\left(\bm{\Omega}\bm{\gamma}(\bm{X})-\bm{\theta}\right)\left(\bm{\Omega}\bm{\gamma}(\bm{X})-\bm{\theta}\right)^{\top}\big{]}\nabla\bm{\eta}(\bm{\theta})^{\top}\bm{\Omega}^{-1}$ (($)$ )
@@
Recalling $\nabla_{\bm{\theta}}u(\bm{\theta}|\bm{\theta}^{(t-1)})=\nabla\bm{\eta}(\bm{\theta})^{\top}\big{(}\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}[\bm{\gamma}(\bm{X})]-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]\big{)},$ which appears in the proof of Proposition 3.1(c), we see that if $\bm{\theta}^{(t)}$ is a stationary point, i.e., $\nabla_{\bm{\theta}}u(\bm{\theta}^{(t)}|\bm{\theta}^{(t-1)})=\bm{0}$, then we have $\mathbb{E}_{p_{\bm{x}}(\bm{\theta}^{(t)})}[\bm{\gamma}(\bm{X})]=\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]$ since $\nabla\bm{\eta}(\theta)$ is non-singular. For $\bm{\theta}=\bm{\theta}^{(t)}$, (15) becomes
$\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t)})}\big{[}\bm{\gamma}(\bm{X})\big{]}$ $=\mathbb{E}_{p_{\bm{x}}(\bm{\theta}^{(t)})}[\bm{\gamma}(\bm{X})]-\big{(}\nabla\bm{\eta}(\bm{\theta}^{(t)})^{\top}\big{)}^{-1}\nabla f(\bm{\theta}^{(t)}).$ (76)
With this and (17), we reach the formula (19) as follows.
@@
$=\bm{\theta}^{(t-1)}-\bm{I}_{p_{\bm{x}}}(\bm{\theta}^{(t-1)})^{-1}\nabla f(\bm{\theta}^{(t-1)}).$ (($)$ )
A.4 Proof of Proposition 4.1
@@
$=\exp\left(-\frac{1}{2}\bm{\theta}^{\top}\bm{\Sigma}^{-1}\bm{\theta}+\frac{1}{2}\big{(}\bm{\Sigma}^{-1}\bm{\theta}-\tilde{\bm{b}}\big{)}^{\top}\big{(}\tilde{\bm{Q}}+\bm{\Sigma}^{-1}\big{)}^{-1}\big{(}\bm{\Sigma}^{-1}\bm{\theta}-\tilde{\bm{b}}\big{)}\right)$
@@
$=\exp\Big{(}-\frac{1}{2}\bm{\theta}^{\top}\bm{Q}\bm{\theta}-\bm{b}^{\top}\bm{\theta}\Big{)}.\qquad(\because\mbox{ definition of }\tilde{\bm{b}})$
Taking the negative logarithm on the both sides leads to (4).
A.5 Proof of Proposition 4.4
@@
$\mathbb{E}_{p_{x}(\theta)}[X(X-1)\cdots(X-n+1)]=m(m-1)\cdots(m-n+1)\theta^{n},$ (77)
@@
$\mathbb{E}_{p_{x}(\theta)}[X(X-1)\cdots(X-n+1)]$ $=\sum_{k=0}^{m}k(k-1)\cdots(k-n+1)\binom{m}{k}(1-\theta)^{m-k}$
@@
where the last equality holds because $\sum\limits_{k=n}^{m}\binom{m-n}{k-n}\theta^{k}(1-\theta)^{m-k}=\sum\limits_{l=0}^{m-n}\binom{m-n}{l}\theta^{l+n}(1-\theta)^{m-(l+n)}=\theta^{n}\sum\limits_{l=0}^{m-n}\binom{m-n}{l}\theta^{l}(1-\theta)^{(m-n)-l}=\theta^{n}$. (77) implies that $\mathbb{E}_{p_{x}(\theta_{j})}\big{[}g_{i,j}(X_{j})\big{]}=\theta_{j}^{n_{i,j}}$, and that
@@
$\equiv K-F(\mathrm{diag}(\bm{u}-\bm{l})\bm{\theta}+\bm{c}).$$\Box$
A.6 Setting of $K$ for (32)

### v2

@@ 5.2. Quadratic program and linear program
$p_{\bm{x}_{1}}(\bm{\lambda})$ $:=\prod_{j\in[p]}\frac{\lambda_{j}^{x_{1,j}}\exp(-\lambda_{j})}{x_{1,j}!}=\underbrace{\Big(\prod\limits_{j\in[p]}x_{1,j}!\Big)^{-1}}_{=\beta(\bm{x}_{1})}\exp\Big(\underbrace{\sum_{j\in[p]}x_{1,j}\ln\lambda_{j}}_{=\bm{\gamma}(\bm{x}_{1})^{\top}\bm{\eta}(\bm{\lambda})}\underbrace{-\sum_{j\in[p]}\lambda_{j}}_{=-A(\bm{\lambda})}\Big)$
@@
$p_{\bm{x}_{2}}(\bm{\theta})$ $:=\frac{1}{(2\pi)^{\frac{n}{2}}\sqrt{|\bm{\Sigma}|}}\exp\Big(-\frac{1}{2}(\bm{x}_{2}-\bm{\theta})^{\top}\bm{\Sigma}^{-1}(\bm{x}_{2}-\bm{\theta})\Big),$
where $\bm{a}_{j}$ and $\bm{q}_{j}$ are the $j$-th column of $\bm{A}$ and $\bm{Q}$, respectively, i.e., $\bm{A}=(\bm{a}_{1},\ldots,\bm{a}_{p})$ and $\bm{Q}=(\bm{q}_{1},\ldots,\bm{q}_{n})$. It is easy to see that (5.20) is a member of the exponential family.
@@
$G_{1}(\bm{x}_{1})=1,\quad G_{2}(\bm{x}_{2})=\sqrt{\frac{|\bm{\Sigma}|}{|\tilde{\bm{Q}}+\bm{\Sigma}^{-1}|}}\exp\Big(-\frac{1}{2}\bm{x}_{2}^{\top}\tilde{\bm{Q}}\bm{x}_{2}-\tilde{\bm{b}}^{\top}\bm{x}_{2}\Big),$
$\tilde{\bm{Q}}:=\bm{\Sigma}^{-1}(\bm{\Sigma}^{-1}-\bm{Q})^{-1}\bm{\Sigma}^{-1}-\bm{\Sigma}^{-1},\quad\mbox{and}\quad\tilde{\bm{b}}:=\bm{\Sigma}^{-1}(\bm{\Sigma}^{-1}-\bm{Q})^{-1}\bm{b},$
the condition (5.6) for (5.19) is fulfilled with $p_{\bm{x}}(\bm{\theta})$ defined by (5.20), $f(\bm{\theta})=\frac{1}{2}\bm{\theta}^{\top}\bm{Q}\bm{\theta}+\bm{b}^{\top}\bm{\theta}$, and $\Theta=\mathbb{R}^{p}_{+}$.
@@
For $\bm{\theta},\bar{\bm{\theta}}\in\mathrm{int}\{\bm{\theta}|\bm{A}^{\top}\bm{\theta}\leq\bm{c}\}$, the surrogate $u(\bm{\theta}|\bar{\bm{\theta}})$ is a convex function of the form:
@@
and the EM algorithm for (5.19) is described in Algorithm 6.
Algorithm 6 EM algorithm for quadratic program (5.19)
1:Let $\bm{\theta}^{(0)}\in\mathbb{R}^{n}$ be an interior point of the feasible region of (5.19), and set $t\leftarrow 1$.
@@
$\bm{\theta}^{(t)}\leftarrow\arg\min_{\bm{\theta}\in\mathbb{R}^{p}}\Big\{\begin{array}[]{r}-\sum\limits_{j\in[p]}\Big(\bm{a}_{j}^{\top}\bm{\theta}+\mu^{(t-1)}_{j}\ln(c_{j}-\bm{a}_{j}^{\top}\bm{\theta})\Big)+\frac{1}{2}\bm{\theta}^{\top}\bm{\Sigma}^{-1}\bm{\theta}+(\bm{\nu}^{(t-1)})^{\top}\bm{\theta}\end{array}\Big\},$ (5.22)
@@
Suppose that $\bm{Q}\in\mathbb{S}^{n}_{+}$, $\bm{A}$ has full row rank, i.e., $\mathrm{rank}A=n$, and (5.19) has an interior point and optimum. Then the sequence, $\{\bm{\theta}^{(t)}\}_{t\geq 0}$, generated by Algorithm 6 converges to an optimal solution $\bm{\theta}^{\star}$ of (5.19). Moreover, if $X^{\star}:=\{\bm{\theta}|f(\bm{\theta})=f(\bm{\theta}^{\star})\}\neq\emptyset$, the sequence $\{\bm{\theta}^{(t)}\}_{t\geq 0}$ converges to an optimal solution of (5.19).
@@
[…]
which is of the form (5.19) with $\bm{c}=(\bm{u}^{\top},-\bm{l}^{\top})^{\top}$ and $\bm{A}=(\bm{I},~-\bm{I})$. For $\bm{\theta}^{(t-1)}=\bar{\bm{\theta}}$, the $\bm{\theta}^{(t)}$ is obtained as the solution $\bm{\theta}$ of
@@
If $\bm{\Sigma}$ is a diagonal matrix $\mathrm{diag}(\bm{\sigma})$ for $\sigma_{j}>0,~j\in[n]$, the $j$-th component of $\bm{\theta}^{(t)}$ is obtained as a solution satisfying $\theta_{j}\in(l_{j},u_{j})$ and
$-\frac{\bar{\theta}_{j}-l_{j}}{\theta_{j}-l_{j}}+\frac{u_{j}-\bar{\theta}_{j}}{u_{j}-\theta_{j}}-\kappa_{j}+\sigma_{j}^{-1}\theta_{j}=0,$ (5.26)
where $\bar{\theta}_{j}=\theta^{(t-1)}_{j}$ and $\kappa_{j}:=[(\bm{\Sigma}^{-1}-\bm{Q})\bar{\bm{\theta}}-\bm{b}]_{j},~j\in[n]$. Note that (5.26) has a unique solution since the left-hand side increases monotonically from -$\infty$ to $\infty$ on the interval $(l_{j},u_{j})$, and it is easy to see that the solution is obtained via the explicit solution formula for the cubic equation:
@@
which is obtained by multiplying (5.26) by $\sigma_{j}(\theta_{j}-l_{j})(u_{j}-\theta_{j})$. Accordingly, the MM algorithm results in a simple component-wise formula like (4.19) and (4.30). (5.25) can also be approached using (4.19), but it is interesting to note that this one is different from (4.19).
@@
We can construct another EM/MM algorithm based on the combination of the Poisson and normal distributions for another form of QP, which is explained in Appendix B since the construction is parallel to the one for (5.19). On the other hand, using this derivation, the EM/MM algorithm for LP can be constructed based only on the Poisson and does not require the normal distribution. See the appendix for the details.
6. Concluding remarks
@@
In particular, while we collectively present several linkages between optimization problems and probability distributions, some problems can be approached via different reformulations and distributions. For example, the quadratic minimization over the rectangle can be solved not only via the approach (ii) in Table 1.1 but also (iv) and (vi), as demonstrated in the final paragraph in the previous section. Further comparison among different formulations or that among different distributions are also to be investigated.
Relations to information geometry may be an interesting research topic as well. The relation to interior point algorithms has been occasionally pointed out in the literature [34, 29, 20], but the analysis based on the exponential family distributions is not examined yet except in the context of statisitical EM algorithms (e.g., [1, 10]). In information geometry, the EM algorithm is understood as optimization on a manifold based on data (e.g., [2]). While a geometric interpretation of the EM algorithm for some estimation problems with probabilistic models has been studied (e.g., [3] for learning hidden variable models in neural networks), it remains to be seen whether these existing perspectives, in conjunction with the EM algorithm developed in this paper, can help to discover new insights into general (non-statistical) optimization problems.
Acknowledgment
The second author’s research is supported in part by MEXT Grant-in-Aid 24K01113. The careful reading and constructive comments of two reviewers have greatly contributed to improving the readability of this paper. Needless to say, any unintentional errors remaining in the final draft are the responsibility of the authors.
Appendix A Derivation of formulae and proof of propositions
A.1. Proof of Proposition 3.1.
(a): The claim is confirmed by simply substituting $\tilde{A}(\bm{\theta}^{(t-1)})$ and $\tilde{\beta}(\bm{x})$ into the formula to be proved.
(b): By the definition (2.5) of the surrogate function $u$, we have
$u(\bm{\theta}|\bm{\theta}^{(t-1)})$
$=-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\ln G(\bm{X})]-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\ln p_{\bm{X}}(\bm{\theta})]+\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\ln q_{\bm{X}}(\bm{\theta}^{(t-1)})]$
$=-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\ln G(\bm{X})]-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\ln p_{\bm{X}}(\bm{\theta})]$
@@ Appendix A Derivation of formulae and proof of propositions
$=-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\ln p_{\bm{X}}(\bm{\theta})]+\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\ln p_{\bm{X}}(\bm{\theta}^{(t-1)})]+f(\bm{\theta}^{(t-1)})$
$=-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]^{\top}\big(\bm{\eta}(\bm{\theta})-\bm{\eta}(\bm{\theta}^{(t-1)})\big)+A(\bm{\theta})-A(\bm{\theta}^{(t-1)})+f(\bm{\theta}^{(t-1)}),$ (3.4)
which shows the formula (3.4).
(c): Although the statement (c) is known in the context of EM algorithms, we here present its proof to make the article self-contained. Note
@@
$=\nabla\bm{\eta}(\bm{\theta})^{\top}\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}[\bm{\gamma}(\bm{X})],$ (A.1)
where all the gradient operations $\nabla$ apply with respect to $\bm{\theta}$. Owing to (3.4) and (A.1), we have
[…]
@@
(d): Observe first that (A.1) holds true also for $\tilde{A}(\bm{\theta})$ and $q_{\bm{x}}(\bm{\theta})$ in place of $A(\bm{\theta})$ and $p_{\bm{x}}(\bm{\theta})$, respectively, i.e.,
$\nabla\tilde{A}(\bm{\theta})=\nabla\bm{\eta}(\bm{\theta})^{\top}\mathbb{E}_{q_{\bm{x}}(\bm{\theta})}[\bm{\gamma}(\bm{X})].$ (A.2)
@@
which leads to (3.5) by multiplying the inverse of $\nabla\bm{\eta}(\bm{\theta})^{\top}$.
A.2. Proof of Corollary 3.2
@@
which equals 0 at $\bm{\theta}=\bar{\bm{\theta}}$.
A.3. Proof of Theorem 3.3
@@
[…]
@@
[…]
@@
$=\bm{I}$ (A.4)
since we have $\int_{\bm{x}\in\Xi}\bm{\Omega}\bm{\gamma}(\bm{x})\nabla p_{\bm{x}}(\bm{\theta})^{\top}\,\mathrm{d}\bm{x}=\bm{I}$ and $\int_{\bm{x}\in\Xi}\nabla p_{\bm{x}}(\bm{\theta})\,\mathrm{d}\bm{x}=\bm{0}$ by differentiating (3.7) and $\int_{\bm{x}\in\Xi}p_{\bm{x}}(\bm{\theta})\,\mathrm{d}\bm{x}=1$ with respect to $\bm{\theta}$, respectively. Therefore, we have
@@
[…]
@@
so we have $\mathbb{V}_{p_{\bm{x}}(\bm{\theta})}[\bm{\Omega}\bm{\gamma}(\bm{X})]=\big(\nabla\bm{\eta}(\bm{\theta})^{\top}\bm{\Omega}^{-1}\big)^{-1}$, establishing the second equality of (3.8). With this, the first equality of (3.8) can also be confirmed as
[…]
@@
Recalling $\nabla_{\bm{\theta}}u(\bm{\theta}|\bm{\theta}^{(t-1)})=\nabla\bm{\eta}(\bm{\theta})^{\top}\big(\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}[\bm{\gamma}(\bm{X})]-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]\big),$ which appears in the proof of Proposition 3.1(c), we see that if $\bm{\theta}^{(t)}$ is a stationary point, i.e., $\nabla_{\bm{\theta}}u(\bm{\theta}^{(t)}|\bm{\theta}^{(t-1)})=\bm{0}$, then we have $\mathbb{E}_{p_{\bm{x}}(\bm{\theta}^{(t)})}[\bm{\gamma}(\bm{X})]=\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]$ since $\nabla\bm{\eta}(\bm{\theta})$ is non-singular. For $\bm{\theta}=\bm{\theta}^{(t)}$, (3.5) becomes
$\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t)})}\big[\bm{\gamma}(\bm{X})\big]$ $=\mathbb{E}_{p_{\bm{x}}(\bm{\theta}^{(t)})}[\bm{\gamma}(\bm{X})]-\big(\nabla\bm{\eta}(\bm{\theta}^{(t)})^{\top}\big)^{-1}\nabla f(\bm{\theta}^{(t)}).$ (A.5)
With this and (3.7), we reach the formula (3.9) as follows.
@@
[…]
A.4. Proof of Proposition 4.1
@@
[…]
@@
[…]
Taking the negative logarithm on the both sides leads to (2.4).
A.5. Proof of Proposition 4.4
@@
$\mathbb{E}_{p_{x}(\theta)}[X(X-1)\cdots(X-n+1)]=m(m-1)\cdots(m-n+1)\theta^{n},$ (A.6)
@@
$\mathbb{E}_{p_{x}(\theta)}[X(X-1)\cdots(X-n+1)]$ $=\sum_{k=0}^{m}k(k-1)\cdots(k-n+1)\binom{m}{k}\theta^{k}(1-\theta)^{m-k}$
@@
[…]
@@
$=K-F(\mathrm{diag}(\bm{u}-\bm{l})\bm{\theta}+\bm{l}).$ $\Box$
A.6. Setting of $K$ for (4.16)